OpenStudy (anonymous):

Find f. Use C for the constant of the first derivative and D for the constant of the second derivative. f '' (x)=24x^3 -18x^2+ 8x

OpenStudy (loser66):

is it not take integral 2 times?

OpenStudy (anonymous):

i think you have to take the anti-derivative twice

OpenStudy (loser66):

the same! antiderivative = integral

OpenStudy (anonymous):

got it still very cofused can either of you help

OpenStudy (anonymous):

would you mind trying the problem

OpenStudy (anonymous):

Using Mathematica 8, the solution to the following differential equation\[f''(x)=24 x^3-18 x^2+8 x \]is\[f(x)=c_2 x+c_1+\frac{6 x^5}{5}-\frac{3 x^4}{2}+\frac{4 x^3}{3} \]

OpenStudy (anonymous):

i got 6x^5/5 -3x^4/2 +4x^2 +Cx+D

OpenStudy (espex):

@alexthomas it looks like you didn't take the integral of your \(4x^2\) term for the second time to get the \(\frac{4}{3}x^3\) before your Cx+D.

OpenStudy (anonymous):

\[f"\left( x \right)=24x ^{3}-18x ^{2}+8x\] \[f \prime \left( x \right)=\frac{ 24x ^{4} }{ 4 }-\frac{ 18x ^{3} }{ 3 }+\frac{ 8x ^{2} }{ 2 }+c\] f'(x)=6x^4-6x^3+4x^2+c similarly you can find f(x)